OpenStudy (anonymous):

A water bottle holds Fraction 5 over 8 liter of water when it is Fraction 3 over 4 full. Which statement best describes the quotient of 5 over 8division sign3 over 4? The maximum amount of water the bottle can hold is Fraction 6 over 5 liters. The amount of water that can be still poured in the bottle is Fraction 5 over 6 liter. The amount of water that can be still poured in the bottle is Fraction 6 over 5 liters. The maximum amount of water that the bottle can hold is Fraction 5 over 6 liter.

OpenStudy (danjs):

\[\frac{ 5 }{ 8 }~Liter =\frac{ 3 }{ 4 }~bottle\]

OpenStudy (danjs):

if you divide both sides by 3/4, you get 1 bottle = #Liters

OpenStudy (danjs):

5/6